# Evaluating and Improving the Robustness of LiDAR-based Localization and Mapping

### v1

@@ Abstract
LiDAR is one of the most commonly adopted sensors for simultaneous localization and mapping (SLAM) and map-based global localization. SLAM and map-based localization are crucial for the independent operation of autonomous systems, especially when external signals such as GNSS are unavailable or unreliable. While state-of-the-art (SOTA) LiDAR SLAM systems could achieve 0.5% (i.e., 0.5m per 100m) of errors and map-based localization could achieve centimeter-level global localization, it is still unclear how robust they are under various common LiDAR data corruptions.
In this work, we extensively evaluated five SOTA LiDAR-based localization systems under 18 common scene-level LiDAR point cloud data (PCD) corruptions. We found that the robustness of LiDAR-based localization varies significantly depending on the category. For SLAM, hand-crafted methods are in general robust against most types of corruption, while being extremely vulnerable (up to +80% errors) to a specific corruption. Learning-based methods are vulnerable to most types of corruptions. For map-based global localization, we found that the SOTA is resistant to all applied corruptions. Finally, we found that simple Bilateral Filter denoising effectively eliminates noise-based corruption but is not helpful in density-based corruption. Re-training is more effective in defending learning-based SLAM against all types of corruption.
@@ I Introduction
Robotics and automation systems have been gaining significant prevalence in this decade. Of those, robots, drones, and autonomous vehicles have been making the news [1] due to their widespread adoption and related safety concerns [2].
Localization is fundamental to these systems since it keeps track of the position of a robotic system within an environment. Outdoors, and Global Navigation Satellite Systems (GNSS) are widely used for localization. In contrast, in urban, tunnel, and indoor settings where GNSS is hindered by weak satellite signals [3], simultaneous localization and mapping (SLAM) systems are favored since they operate independently of external signals and reliably navigate both indoors and outdoors. For these reasons, SLAM is widely used for locating and mapping-related tasks in robotics and autonomous driving. SLAM methods utilize either a camera, LiDAR, or a combination of both, with the optional integration of an Inertial Measurement Unit (IMU) [4]. Compared to vision SLAM, LiDAR SLAM has higher mapping accuracy, and better stability, is less susceptible to lighting conditions, and does not observe scale drift. These advantages lead to its popularity in large-scale use cases and facilitate more research and application in this direction.
Despite the recent successes of LiDAR SLAM, many challenges remain. Laser signals are robust under clear conditions, however, they decay significantly in rainy and foggy weather [5], making LiDAR SLAM degrade in adverse weather conditions [4]. Noise commonly occurs in point cloud data due to hardware limitations, motion, lighting, dust, or reflective surface [6, 7, 8, 9, 10, 11]. Also, mechanical, solid-state, and phased-array LiDAR sensors are built differently and SLAM designed for one may not be interchangeable with others [4]. Instability causes localization to miscalculate a deviation from the agent’s current position. Depending on the scale of the deviation, it can cause automated vehicles to drive off-road or onto the wrong lane. In robotic systems, it would result in the wake-up or kidnapped robot problem, since the large gap between the two consecutive frames’ positions can cause the robot to be unable to recover, i.e., either it stops due to failed localization or reset and initiates a costly recovery and relocalize process, discarding the learned map/knowledge [12, 13, 14].
Such challenges to the LiDAR sensor raise concerns about the performance of SLAM in diverse environments [4]. To this end, there is a lack of comprehensive evaluation on the manifestation and impact of such issues on LiDAR SLAM. Moreover, increasing efforts invested in learning-based SLAM raises additional robustness concerns due to DNNs’ sensitivity to noises [15, 16, 17]. Yet there is a lack of comparative evaluation in that regard. Hence in this work, we evaluate the robustness of such applications against common LiDAR data corruption due to adverse weather, noisy environments, and losses. Once the effects are discovered, we want to explore whether current defense strategies such as re-training and denoising are sufficient to address these issues.
Our contributions can be summarized as follows:
  • We propose a novel framework to evaluate the robustness of the SOTA LiDAR-based localization against common LiDAR data corruptions. To the best of our knowledge, this is the first work to study this problem.
  • We study the effectiveness of common defense strategies, such as denoising and re-training, against common LiDAR data corruption.
@@
  • We share our code and data publicly [^1] https://github.com/boyang9602/LiDARLocRobustness.
@@ II Background and related works
LiDAR-based localization can either simultaneously build a map and localize the agent within it, i.e., SLAM [4], or localize the agent within a pre-built map, i.e., global localization [18].
II-A LiDAR-based localization and mapping
Figure 1: Different types of LiDAR SLAM Overview.
LiDAR-based localization is the most accurate, stable, resistant to different lighting conditions, and has no scale drift, thus suitable for large-scale applications [4]. It could be used from autonomous driving [19, 20], robot exploration [21] to deep space exploration [22]. LiDAR-based localization can be roughly categorized into LiDAR-only SLAM, MSF-based SLAM, and map-based global localization.
II-A1 LiDAR-only SLAM
LiDAR-only SLAM can be categorized into feature-based, direct, and projection-based methods as shown in Figure 1. Feature-based methods [23, 24] extract appearance, geometric, semantics, or shape-based features for scan matching. Alternatively, direct methods [25, 26] match the scan directly via handcrafted point cloud registration methods, neural networks, or a map. Finally, projection-based approaches [27, 28] convert the point cloud to a 2D representation and leverage existing visual SLAM methods. The effectiveness of these methods is highly dependent on the environments, equipment, and platforms. Traditional SLAM frameworks use hand-crafted features or models [29, 30, 31, 32, 33] to register scans. In contrast, learning-based SLAM [34, 35, 25], facilitated by the development of 3D computer vision, extracts features and register scans automatically. Due to the focus on computer vision, these methods only focus on feature extraction and scan registration rather than a complete solution. Moreover, their accuracy, stability, and performance leave much to be desired [4]. In our work, we look to evaluate the robustness of LiDAR-only SLAM.
II-A2 MSF-based SLAM
On the other hand, LiDAR can be incorporated into an MSF-based SLAM. MSF SLAM combines the output from the LiDAR with camera, IMU, and GNSS, to complement each other and improve localization performance [4, 36]. These methods build upon the LiDAR-only solutions. They are not included in this study.
II-A3 Map-based LiDAR Localization
Map-based localization registers an online LiDAR scan with a pre-built map and tries to locate it within the map [18, 37, 38, 39, 40, 41]. This offers precise and efficient localization, which is especially useful in safety-critical systems such as autonomous driving systems [42, 43]. However, these methods require a pre-built map, which is typically created using the SLAM algorithm offline and may require ground truth information to improve the map quality. Map-based localization shares many common technologies with LiDAR-only SLAM (e.g., LiDAR scan registration), so we will also attempt to evaluate the robustness of such systems.
@@ II-B Robustness of LiDAR-based Systems
Previous works evaluated the impact of data corruption on LiDAR-based 3D obstacle detection. This includes benchmarking against perturbations [44, 45, 46, 47, 48]; testing and adversarial attacks via point-based [49, 50, 51, 52, 53, 54, 55, 56, 57] and object spoofing methods [58, 59, 60, 61, 62, 63, 64]. Finally, complex weather simulators [65, 66, 67, 68] model LiDAR’s interaction with dense droplets as well as how light bounces off objects under rain, snow, and fog. Many of these works are tailored to objects and thus inapplicable to localization. Hence, in this work, we evaluate LiDAR-based localization against a set of applicable perturbations.
@@ III Methodology
In this section, we describe our framework for evaluating and improving the robustness of LiDAR-based localization.
@@
III-A Evaluating the Robustness of LiDAR SLAM
LiDAR SLAMs are deployed in complex and diverse environments which may impact the sensitivity of LiDAR. It is important to evaluate the robustness and reliability of such systems in various environments. Our framework employs a total of 18 perturbations on cloud corruptions to emulate diverse environments such as adverse weather conditions and equipment malfunctions. We select a comprehensive set of corruption reflecting commonly occurring phenomena that impact the entire LiDAR scan instead of individual objects. Table I summarizes the list of perturbation types.
Weather Perturbations. Adverse weather such as rain and snow can negatively affect the quality of LiDAR scans. Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, thus injecting noises into the resulting PCD. Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which will reflect differently from dry ground.
Figure 2: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data.
@@ III-A Evaluating the Robustness of LiDAR SLAM
  • Rain or Snow. We utilize an off-the-shelf weather simulator, i.e., LISA [65, 66] for simulating PCD under adverse weather such as rain and snow. LISA utilizes a physics-based model to simulate the decayed return of laser pulses in rain and snow weather, which is realistic as evaluated by previous work [47].
  • Fog. We utilize a fog Simulation on Real LiDAR Point Clouds (FS) [67] to simulate fog. While LISA already has a fog perturbation, it is more general, and works such as FS provide a better fog simulation model without needing unavailable data such as optical channel information. Hence, we use FS for fog simulation. FS simulates the effect of fog on real-world LiDAR point clouds that have been recorded in clear weather. FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
  • Rain + Wet Ground (rain_wg). We combine LISA and Wet Ground Model from [68] to produce a more realistic simulation than rain alone. The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water. Figure 2 shows an example of this corruption.
  • Snow + Wet Ground (snow_wg). Similar to rain + wet ground, we produce more realistic snow weather by combining LISA’s snow simulation and the wet ground model on PCD.
Noise Perturbations. Noises happen during LiDAR measurement due to internal (e.g., equipment vibration and measurement errors) or external (e.g., dust and background light) factors [11, 69, 70]. We design perturbations on PCDs to simulate the errors attributed to such internal and external factors.
@@
TABLE II: Noise-based corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. RS refers to a random subset function, which samples a subset from the input based on the given $SEV$. $\Delta p$ refers to a point in CCS, so $+\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $+\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of PCD is annotated as $P=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to emulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [48] while SCS can better represent inaccurate measurement of time of flight (ToF) [47].
In addition to re-positioning existing PCD points, we synthesize new points to emulate environment noises, such as dust in the air, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
@@
TABLE III: Density-based corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. IP refers to the 3D interpolation function. $NN_{k}$ refers to the k-nearest neighbor function. RS refers to a random subset function, which samples a subset from the input based on the given number.
@@
III-B Improving the Robustness of LiDAR SLAM
Re-training. Learning-based systems are data-driven and could learn to fit themselves into a data distribution. We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning [65, 67, 68, 73, 74].
Denoising. Handcraft algorithms are expected to be tuned and adapted by experts once inefficiencies are identified. Therefore such improvements are achieved case by case and lack of generality. Hence in this work, for non-learning-based SLAM, we focus on denoising the data, which is a more general way. Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality. There are three categories of denoising, i.e., filter, optimization, and deep learning-based, each with its advantages and disadvantages [11]. In this work, we adopted a filter-based method, i.e., the bilateral filter by [75], which is reliable, and significantly faster than the alternatives while preserving the intensity with readily available implementation.

### v2

# Robustness of LiDAR-Based Pose Estimation: Evaluating and Improving Odometry and Localization Under Common Point Cloud Corruptions

@@ Abstract
Accurate and reliable pose estimation, i.e., determining the precise position and orientation of autonomous robots and vehicles, is critical for tasks like navigation and mapping. LiDAR is a widely used sensor for pose estimation, with odometry and localization being two primary tasks. LiDAR odometry estimates the relative motion between consecutive scans, while LiDAR localization aligns real-time scans with a pre-recorded map to obtain a global pose. Although they have different objectives and application scenarios, both rely on point cloud registration as the underlying technique and face shared challenges of data corruption caused by adverse conditions (e.g., rain). While state-of-the-art (SOTA) pose estimation systems achieved high accuracy on clean data, their robustness to corrupted data remains unclear. In this work, we propose a framework to systematically evaluate five SOTA LiDAR pose estimation systems across 18 synthetic real-world point cloud corruptions. Our experiments reveal that odometry systems degrade significantly under specific corruptions, with relative position errors increasing from 0.5% to more than 80%, while localization systems remain highly robust. We further demonstrate that denoising techniques can effectively mitigate the adverse effects of noise-induced corruptions, and re-training learning-based systems with corrupted data significantly enhances the robustness against various corruption types.
@@ I Introduction
Autonomous robots and vehicles rely on real-time pose estimation to navigate, avoid collisions, and perform tasks such as mapping and path planning. LiDAR enables precise environmental perception by generating high-resolution 3D point clouds independent of external lighting conditions. As a result, it is widely used for pose estimation, particularly in scenarios where GNSS signals are unreliable (e.g., urban canyons) or unavailable (e.g., tunnels). Two primary methods are commonly employed for LiDAR-based pose estimation: odometry [1] and localization [2]. LiDAR odometry estimates incremental motion by registering consecutive scans, making it essential for exploring unknown environments. However, odometry suffers from drift due to error accumulation over time. LiDAR localization, on the other hand, aligns online scans with pre-built maps to achieve drift-free global pose estimation. This is particularly crucial in safety-critical applications, such as fully autonomous vehicles that require global localization accuracy at the centimeter level [3]. Despite the different objectives, both tasks share underlying point cloud registration techniques to obtain fine poses.
Point cloud data corruptions can occur in various real-world scenarios, including adverse weather [4], noises from sensor [5] or dusts [6], reflective surfaces [7], and occlusions caused by dynamic environments [8, 9]. Despite SOTA LiDAR-based odometry and localization methods achieving highly accurate results on clean LiDAR datasets, corrupted point cloud data will inevitably bring challenges to their performance [10]. However, the extent to which their performance degrades under these corruptions remains unclear. Severe performance degradation could lead to critical failures, such as robots getting lost [11] and autonomous vehicles driving off-road [12]. Though prior studies explored the robustness [9] and security [13] of LiDAR-based localization systems, there is still a lack of systematic evaluation of how point cloud corruptions impact LiDAR-based pose estimation systems. Our work bridges this gap with three key contributions:
  • A framework to systematically evaluate the robustness of five SOTA LiDAR-based pose estimation systems under 18 common point cloud corruptions.
  • An investigation of mitigation strategies, including denoising and re-training (for learning-based systems).
@@
  • Public release of code and data for reproducibility [^1] https://github.com/boyang9602/LiDARLocRobustness.
@@ II Background and related works
II-A LiDAR-Based Pose Estimation
LiDAR-based pose estimation allows autonomous agents to track their motions (odometry) and get their absolute position and orientation in a map (localization) in 3D environments. The core of LiDAR-based pose estimation is point cloud registration, i.e., aligning LiDAR scans to estimate their 6-degree-of-freedom (DoF) spatial transformations (translation and rotation).
II-A1 Point Cloud Registration
Given a source point cloud $\mathcal{P}\in\mathbb{R}^{3\times N}$ and a target point cloud $\mathcal{T}\in\mathbb{R}^{3\times M}$, point cloud registration solves:
$\mathbf{R}^{*},\mathbf{t}^{*}=\underset{\mathbf{R}\in SO(3),\mathbf{t\in\mathbb{R}^{3}}}{\arg\min}\ \text{dist}\left(\mathbf{R}\mathcal{P}+\mathbf{t},\mathcal{T}\right)$ (1)
Here, $\mathbf{R}$ is a 3-DoF rotation matrix, and $\mathbf{t}$ is a 3-DoF translation vector. The operation $\mathbf{R}\mathcal{P}+\mathbf{t}$ rotates all points using $\mathbf{R}$ and translates them by $\mathbf{t}$. The function $\text{dist}(\cdot,\cdot)$ quantifies the errors between the corresponding points in $\mathbf{R}\mathcal{P}+\mathbf{t}$ and $\mathcal{T}$. In other words, this equation seeks the optimal pose $\mathbf{R}^{*},\mathbf{t}^{*}$ that aligns the source point cloud $\mathcal{P}$ to the target point cloud $\mathcal{T}$.
Various solutions have been proposed for point cloud registration, evolving from geometry-based methods to learning-based methods. They can be broadly categorized into direct matching, feature-based, and learning-based methods [1]. Direct matching methods, such as Iterative Closest Point (ICP) [14], align raw points by iteratively minimizing the distance between nearest neighbor points. While vanilla ICP struggles with problems such as sensitivity to initialization and noise, modern variants such as KISS-ICP [15] address these limitations and remain competitive currently. Another direct method, Normal Distribution Transform (NDT) [16], partitions the target point cloud into a grid of cells, with each cell modeling the local distribution of points as a Gaussian. Then it finds the optimal pose that transforms the source point cloud to best fit the normal distribution in the target grid. Feature-based methods like LiDAR Odometry and Mapping (LOAM) [17] achieve fast and robust odometry by extracting and matching feature points such as sharp edges and planar surfaces. With the development of deep learning, researchers further improve feature extraction from hand-crafted features to learned features and predict the pose using neural networks [18]. Instead of processing the 3D point cloud directly, [19] projects the points into 2D images and leverages the neural network to extract features from images and predict the pose. The success of Neural Radiance Fields (NeRF) [20] further inspires the recent exploration of implicit neural representations of scenes and advances the LiDAR-based pose estimations [21, 22].
II-A2 LiDAR Odometry
LiDAR odometry tracks the accurate motion of autonomous agents in real-time by aligning consecutive LiDAR scans. Despite the high accuracy of point cloud registration, the long-term trajectory drifts significantly due to accumulated errors. To mitigate the drift, additional algorithms such as loop closure [23] should be applied in post-processing. SOTA odometry systems cover all the approaches mentioned in the previous section [15, 17, 24, 19, 18, 22].
II-A3 LiDAR Localization
LiDAR localization determines the global 6-DoF pose by aligning an online scan with a pre-built map, which is critical for fully autonomous driving systems that require centimeter-level accuracy. While some localization systems estimate only 3-DoF position, we focus on 6-DoF pose estimation in this work. Compared to scan-to-scan registration in odometry, scan-to-map registration in localization is more challenging due to the vast search space. To address this problem, most methods first obtain an initial pose via external signals (e.g., GNSS) [3] or global position recognition [2], then perform the point cloud registration to estimate the fine pose. Although one-stage global pose estimation approaches exist, they generally underperform compared to those with initialization [2]. Similar to odometry, SOTA localization approaches also cover the full spectrum of registration methods [3, 25, 26, 21].
II-A4 Multi-Sensor Fusion Approaches
Multi-sensor fusion (MSF) combines complementary sensors, such as LiDAR, cameras (vision odometry and localization), IMUs (high-frequency motion data), and GNSS (global localization), to enhance pose estimation [3]. In this work, we focus on LiDAR-only pose estimation systems to study LiDAR’s standalone vulnerabilities in pose estimation.
@@ II-B Robustness of LiDAR-based Systems
Previous research has extensively explored the robustness and security of LiDAR-based 3D perception systems, particularly in obstacle detection. Several studies have proposed realistic adverse weather simulations for point cloud, including rain [27, 28], snow [29], and fog [30]. Researchers also systematically benchmarked the LiDAR-based object detection on common point cloud data corruptions [31] and investigated its vulnerability to adversarial attacks [32].
In the domain of LiDAR-based pose estimation, recent works have examined the security of localization and odometry against adversarial attacks [13, 33, 34]. The most relevant study to ours is [9], which proposed a closed-form approximation of localization errors for the ICP algorithm. In contrast, our study systematically evaluates the robustness of multiple pose estimation systems, covering various underlying point cloud registration algorithms, against a comprehensive set of point cloud data corruptions.
@@ III Methodology
This section presents our framework for evaluating and improving the robustness of LiDAR-based pose estimation.
@@
III-A Evaluating Robustness of LiDAR-Based Pose Estimation
Autonomous robotics and vehicles are deployed in complex and diverse environments which affects the LiDAR performance and, consequently, the accuracy of pose estimation. To rigorously evaluate the robustness of pose estimation systems, our framework employs a broad spectrum of point cloud corruption (18 different types) that simulates adverse weather, internal and external noises, equipment malfunctions, and occlusions. Table I summarizes the list of corruption types.
Weather Perturbations. Adverse weather can negatively affect the quality of LiDAR scans. Randomly distributed droplets in the air from rain, snow, and fog can scatter light and disturb the measurement of LiDAR, and cause corruptions in point cloud. Furthermore, wet ground always co-occurs with rain and snow due to the precipitation, which reflects differently from dry ground. Despite the availability of real-world datasets capturing various adverse weather conditions [35, 36], we opt for simulation methods on clean datasets for the following reasons:
  • Realism. Existing simulation methods are physics model-based and highly resemble real-world phenomena [31, 27, 29, 30].
  • Controllable. In real-world datasets, the severity of adverse weather is typically unknown and cannot be adjusted. Simulation, on the other hand, allows us to parameterize the intensity of these conditions.
  • Sensor limitations. Real-world datasets are limited by their specific sensors, which may be incompatible with certain algorithms [10]. Simulation allows us to apply adverse weather on any dataset.
  • Completeness. No dataset contains the full range of corruptions. Simulation allows us to apply all corruptions on a single dataset for a comprehensive and fair comparison.
Figure 1: An example of rain and wet ground corruption. The field view and reflection intensities are significantly decayed compared to the clean data. The color bar indicates the intensities of the points.
@@ III-A Evaluating Robustness of LiDAR-Based Pose Estimation
  • Rain or Snow. We utilize an off-the-shelf weather simulator, i.e., LISA [27, 28] to simulate rain and snow. LISA utilizes a physics model to simulate the decayed return of laser pulses in rain and snow weather.
  • Fog. We utilize Fog Simulation (FS) [30] to simulate fog. FS models fog by modifying the impulse response changes to reflect the impact of different fog densities on the environment and the surrounding objects.
  • Rain or Snow + Wet Ground (rain/snow_wg). We combine LISA and Wet Ground Model from [29] to produce a more realistic simulation than rain/snow alone. The Wet Ground Model builds an optical model to simulate the reflection on the wet ground based on the thickness of the water. Figure 1 shows an example of this corruption.
Noise Perturbations. Noises occur in LiDAR measurement due to internal (e.g., equipment vibration) or external (e.g., dust) factors [37, 38]. We design perturbations on point cloud to simulate the errors attributed to such factors.
@@
TABLE II: Noise-induced corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. RS refers to a random subset function, which samples a subset from the input based on the given $SEV$. $\Delta p$ refers to a point in CCS, so $+\Delta p$ means we change the three coordinates in CCS. $\Delta r$ refers to the range value, $+\Delta r$ means we only change the range value in SCS.
We show the details of noise perturbations in Table II. One frame of point cloud is annotated as $\mathcal{P}=\{p_{i}|i=1,2,...N\}$, where $N$ is the number of points in the point cloud. We apply three distributions of noises (Gaussian, Uniform, and Impulse) to two coordinate systems (Cartesian and Spherical, CCS and SCS for short) to simulate different noise sources. In CCS, one point is represented by $p=\{x,y,z\}$, i.e., Euclidean distance to each planes, while in SCS, one point is represented by $p=\{r,\phi,\theta\}$, where $r$ is the range, i.e., Euclidean distance from the origin to the point, $\phi$ is the azimuthal angle and $\theta$ is the polar angle. Noises in CCS represent inaccuracies caused by the vibration or rotating of the sensor [39] while SCS can better represent inaccurate measurement of time of flight (ToF) [31]. In addition to re-positioning existing LiDAR points, we synthesize new points to simulate environment noises, such as dust in the air, through background noise and upsample applied to CCS. Background noise adds points randomly in the entire space, while upsample adds points close to randomly selected points.
@@
TABLE III: Density-related corruptions. SEV (severity) refers to the significance of corruption, whose definition varies based on the type of corruption. C refers to a constant value, whose concrete value varies in different places. IP refers to the 3D interpolation function. $NN_{k}$ refers to the k-nearest neighbor function. RS refers to a random subset function, which samples a subset from the input based on the given number.
@@
III-B Improving Robustness of LiDAR-Based Pose Estimation
Re-training. Learning-based systems are data-driven and could learn to fit themselves into a data distribution. We expect these systems to learn how to handle the corrupted data by themselves through re-training, which is a common enhancement and defensive strategy for deep learning-based LiDAR-based algorithms [27, 29, 30].
Denoising. Handcraft algorithms are expected to be tuned and adapted by experts once inefficiencies are identified. Therefore such improvements are achieved case by case and lack of generality. Hence in this work, for non-learning-based pose estimation systems, we focus on denoising the data, which is a more general way. Point cloud denoising removes unwanted noises from LiDAR scans to improve their quality. There are three categories of denoising, i.e., filter, optimization, and deep learning-based, each with its advantages and disadvantages [37]. In this work, we adopted a filter-based method, i.e., the bilateral filter by [40], which is reliable, and significantly faster than the alternatives while preserving the intensity with readily available implementation.